# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 8. Proof of main theorems
The constant appearing is
$C_{8}=50000\pi.$
First notice that
$\|a_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|a_{g}\|_{\infty}.$
Since the sequence $(a_{g})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{g}\|\leq 2\,a_{max}.$
Hence the previous limit supremum expression is bounded from above by
$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
Using Theorem 8.2, we have that
$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
$\leq\ \frac{1}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(\frac{3}{16}-\eta\right)^{3}}\,\left(|X_{g}|+8\pi^{2}\,|X_{g}|\,\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
with propability $1-\mathcal{O}(g^{-c(\eta)})$, where $\eta\in(0,3/16)$, $c(\eta)$ is a positive constant depending on $\eta$, and where
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
@@
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+8\pi^{2}\,e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
This means that with high propability Expression (8.4) equals
$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta$
where equality followed from using that $T\,\delta=\pi/2$. $C_{10}$ is a constant upper bound coming from
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}.$
If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
@@ Acknowledgements.
The author would like to thank Prof. Mostafa Sabri, Prof. Tuomas Sahlsten, and Prof. Farrell Brumley for discussions that led to this article. The author also thanks Prof. Tuomas Sahlsten, Prof. Henrik Ueberschär, Dr. Félix Lequen, Dr. Søren Mikkelsen, and Dr. Joe Thomas for their guidance while writing this article.
Our use of the exponential is inspired by the upcoming article [HLM+] of the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten and Henrik Ueberschär on quantum ergodicity for eigenfunctions for Schrödinger operators on hyperbolic surfaces. We thank the other authors of that paper.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University.

### v2

@@ 8. Proof of main theorems
The constant appearing is $C_{8}=5120000\pi^{2}$.
First, notice that
$\|\bar{a}_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|\bar{a}_{g}\|_{\infty}.$
Since the sequence $(a_{g})$ is uniformly bounded, we have
$\|\bar{a}_{g}\|_{\infty}\leq 2\|a_{g}\|_{\infty}\leq 2a_{max},$
where $a_{max}$ is the universal bound for $(a_{g})$. Hence the previous limit supremum expression is bounded from above by
$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
By using Theorem 8.2, we have
$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
$\leq\ \frac{|X_{g}|}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(1-\sqrt{1-4(\tfrac{1}{4}-\eta)}\right)^{3}}\,\left(1+\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
with probability 1o(1)1-o(1), where $\eta\in(0,\tfrac{1}{4})$ and
$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
@@
$\frac{1}{\left(1-\sqrt{4\eta}\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
This means that when passing to the high-genus limit, we have
$\frac{1}{\left(1-\sqrt{4\eta}\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right)\xrightarrow{g\to\infty}C_{10}$
for some constant $C_{10}$. This means that with high probability, Expression (8.4) has the following upper bound:
$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta,$
where the equality follows from using $T\delta=\tfrac{\pi}{2}$.
If we had originally chosen
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\tfrac{2}{9}\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice at the beginning of the proof to conclude the argument.
Appendix A On quantum mixing of physical-space observables on flat 2-tori
Let $\mathbb{T}_{L}^{2}$ be a flat $2$-torus with side length $L$. We consider the sequence $\{\mathbb{T}_{j}^{2}\}_{j\in\mathbb{N}}$ of growing flat $2$-tori obtained by letting $L=j$. The torus $\mathbb{T}_{j}^{2}$ has eigenvalues and corresponding eigenfunctions
$\lambda_{(m,n)}^{(j)}=\frac{4\pi^{2}}{j^{2}}(m^{2}+n^{2}),\qquad\psi_{(m,n)}^{(j)}(x,y)=e^{\frac{2\pi i}{j}(mx+ny)},$
where $(m,n)\in\mathbb{Z}^{2}$. Let $\rho_{(m,n)}^{(j)}=\sqrt{\lambda^{(j)}_{(m,n)}-\tfrac{1}{4}}$. We consider Properties (II) and (III) of Theorem 1.3 and show that the corresponding properties fail for the sequence $\{\mathbb{T}_{j}^{2}\}$. In fact, it suffices to show that Property (II) fails, since this already implies the failure of Property (III).
Proposition A.1 (Large-scale quantum mixing breaks).
Let $I\subset(0,\infty)$ be a compact interval. There exists $\epsilon>0$ and a sequence of bounded functions $\{a_{j}:\mathbb{T}_{j}^{2}\to\mathbb{C}\}$ such that for every $\delta>0$ we have that
$\limsup_{j\to\infty}\frac{1}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\,\sum_{\begin{subarray}{c}(p,q)\in\mathbb{Z}^{2}:\rho_{(p,q)}^{(j)}\in I,(p,q)\neq(m,n),\\ |\rho_{(m,n)}^{(j)}-\rho_{(p,q)}^{(j)}|<\delta\end{subarray}}|\langle\psi_{(m,n)}^{(j)},a\psi_{(p,q)}^{(j)}\rangle|^{2}\geq\epsilon.$
Proof of Proposition A.1.
Let $\epsilon>0$ be specified later, and define $a_{j}(x,y)=e^{\frac{2\pi i}{j}(x+y)}$. Let $\delta>0$ be arbitrary; in particular, it is independent of $\epsilon$ and the choice of $a_{j}$.
We estimate the double sum appearing in the statement from below. A convenient lower bound is obtained by restricting to pairs of indices of the form $(m,n)$ and (m1,n1)(m-1,n-1):
$\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}}\,|\langle\psi_{(m,n)}^{(j)},a_{j}\psi^{(j)}_{(m-1,n-1)}\rangle|^{2}$
$=\ \sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}},$
where the equality follows from the identity
$a_{j}\psi_{(m-1,n-1)}^{(j)}=\psi_{(m,n)}^{(j)}.$
Consequently, the $\limsup$ in the statement admits the lower bound
$\limsup_{j\to\infty}\frac{1}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}}$
$=\ \limsup_{j\to\infty}\frac{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}|<\delta\}}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}.$
As $j\to\infty$, the spacing between $\rho_{(m,n)}^{(j)}$ and $\rho_{(m-1,n-1)}^{(j)}$ tends to zero uniformly for indices with $\rho_{(m,n)}^{(j)}\in I$. In particular, for sufficiently large $j$, the condition $|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta$ is automatically satisfied whenever both terms lie in $I$. Moreover, we have the asymptotic relation
$\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I\}\sim\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}.$
It follows that
$\limsup_{j\to\infty}\frac{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}|<\delta\}}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}=C$
for some constant $C>0$. Choosing $\epsilon=\tfrac{C}{2}$ yields the desired conclusion.
In Subsection 1.1, we noted that it is an interesting problem whether analogues of Properties (II) and (III) of Theorem 1.1 hold for a fixed flat $2$-torus. The example above does not readily extend to that setting: when the torus is fixed and one considers higher and higher eigenvalues, the spectrum does not exhibit the same clustering, and nearby eigenvalues do not accumulate in the same way.
@@ Acknowledgements.
The author would like to thank Prof. Farrell Brumley, Prof. Mostafa Sabri, and Prof. Tuomas Sahlsten for discussions that led to this article. The author also thanks Prof. Tuomas Sahlsten, Prof. Henrik Ueberschär, Dr. Félix Lequen, Dr. Søren Mikkelsen, and Dr. Joe Thomas for their guidance while writing this article. Finally, the author wants to thank Prof. Zeév Rudnick and Prof. Mostafa Sabri for their helpful comments on the earlier drafts.
Our use of the exponential is inspired by the upcoming article [30] by the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten, and Henrik Ueberschär on quantum ergodicity of eigenfunctions for Schrödinger operators on hyperbolic surfaces. We thank the other authors of that paper.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, and the Research Council of Finland, project: Quantum chaos on large and many-body systems, grant numbers 347365, 353738, and Aalto University.